OpenStudy (anonymous):

Divide (-10x^3 +30x -20) / (-5x +5)

OpenStudy (anonymous):

@thomaster

OpenStudy (agent0smith):

You can simplify it a little first by pulling out a factor of 10 on the top, and 5 on the bottom, then try to either factor it, or use long division \[\Large \frac{ 10(-x^3 +3x -2)}{5(-x +1) }\]

OpenStudy (anonymous):

but the x^3 +3x-2 doesnt factor

OpenStudy (agent0smith):

Then you'll have to use either long division, or synthetic division. Do you know how to do either of those?

OpenStudy (anonymous):

no

OpenStudy (agent0smith):

You haven't done any polynomial long division in class??

OpenStudy (anonymous):

im out sick alot. i must have missed it when they went over it

OpenStudy (agent0smith):

Then you might want to watch this, since if i show you how to do it, you might not follow: http://www.youtube.com/watch?v=l6_ghhd7kwQ Or maybe try this site: http://www.mathsisfun.com/algebra/polynomials-division-long.html

OpenStudy (anonymous):

ok

OpenStudy (agent0smith):

Once you simplify the 10/5 and distribute, you get\[\Large \frac{ 10(-x^3 +3x -2)}{5(-x +1) } = \] \[\Large \frac{ (-2x^3 +6x -4)}{(-x +1) }\] Then set up the division like so, putting in a 0x^2 since there's no x^2 term. 

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

2x^2 +2x -4 is the answer i got